# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v2

@@ Abstract
We show that a minor variant of the continuous Fréchet distance between polygonal curves can be computed using essentially the same algorithm used to solve the discrete version, thus dramatically simplifying the algorithm for computing it. The new variant is not necessarily monotone, but this shortcoming can be easily handled via refinement.
Combined with a Dijkstra/Prim type algorithm, this leads to a realization of the Fréchet distance (i.e., a morphing) that is locally optimal (aka locally correct), that is both easy to compute, and in practice, takes near linear time on many inputs. The new morphing has the property that the leash is always as short-as-possible. These matchings are more natural, and are “better” than the ones computed by standard algorithms. This should make the Fréchet distance more attractive for real world applications.
We implemented the new algorithm, and developed various strategies to obtain fast execution in practice. Among our contributions is a new simplification strategy that is distance-sensitive, and enables us to compute the exact continuous Fréchet distance in near linear time in practice. We preformed extensive experiments on our new algorithm, and released Julia and Python packages with these new implementations.
Our implementations are available here (Python) and here (Julia). Animations and examples are available here.
@@ 1.1 Definitions
Given two polygonal curves, their Fréchet distance is the length of a leash that a person needs, if they walk along one of the curves, while a dog connected by the leash walks along the other curve, assuming they synchronize their walks so as to minimize the length of this leash. (I.e. they walk so as to minimize their maximum distance apart over the walk.) Our approach is (slightly) different than the standard approach, and we define it carefully first.
@@ 1.1.2 Fréchet distance
Conceptually, the Fréchet distance is the problem of computing minimum bottleneck matching between the two curves, respecting the order and continuity along the curves. [^2] Formally, since the reparameterization is not one-to-one, this is not quite a matching. One can restrict to using only such bijections, with no adverse effects, but it adds a level of tediousness, which we avoid for the sake of simplicity of exposition. Alternatively, it is an $L_{\infty}$-norm type measure of the similarity between two curves. It thus suffers from sensitivity to outliers. Furthermore, even if only a small portion of the morphing requires a long leash, the measure, and the algorithms computing it, may use this long leash in large portions of the walk, generating a matching that is loose in many places, see Figure 1.2.
@@
Figure 1.1: Two curves, their free space diagram (and the associated elevation function), and the optimal Fréchet morphing between the two curves encoded as an $x/y$-monotone curve.
(a) The classical (discrete) Fréchet morphing, caring only about the maximum leash length.
(b) The retractable discrete Fréchet morphing, using the shortest leash possible at each point.
@@ 1.2 Background
Alarmingly, these algorithms “inherently” require quadratic time, although a logarithmic speedup is possible bbmm-fswdi-17. It is believed the problem requires quadratic time in the worst case (see bbmm-fswdi-17 and references therein). The quadratic time can be improved for realistic inputs by assuming that the input is “nice”, and introducing approximation, but the resulting algorithms are still not simple dhw-afdrc-12.
@@ 1.2.1 Variants of the Fréchet distance
The weak Fréchet distance allows morphings where the agents are allowed to go back and traverse portion of the curve already traveled (i.e., the morphing does not have to be $x/y$-monotone). Since the weak Fréchet distance allows considering more parameterizations, it is potentially smaller, and we have that $\mathsf{d}^{w}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, for any two curves $\pi,\sigma$. Elegantly, Alt and Godau ag-cfdbt-95 showed the weak Fréchet distance can be reduced to computing the minimum spanning tree of the appropriate graph. Unfortunately, there does not seem to be a natural way to overcome this non-monotonicity (and thus get the “strong” version).
@@
The complexity of these algorithms, together with sensitivity of the Fréchet distance to noise, lead to using “easier” related measures, such as the discrete version of the problem, and dynamic time-warping (discussed below). In the discrete version, you are given two sequences of points $p_{1},\ldots,p_{n}$, and $q_{1},\ldots,q_{m}$, and the purpose is for two “frogs” starting at $p_{1}$ and $q_{1}$, respectively, to jump through the points in the sequence until reaching $p_{n},q_{m}$, respectively, while minimizing the maximum distance between the two frogs during this traversal. At each step, only one of the frogs can jump from its current location, to the next point in its sequence (i.e., no jumping back). Computing the optimal distance under this measure can be done by dynamic-programming similar to the standard approach to edit-distance. Indeed, the configuration space here is the grid $H=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$, where $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$.
@@
For the (monotone) discrete Fréchet distance the induced graph on the grid $H$ is a DAG, and the task of computing the Fréchet distance, is to find a minimum bottleneck path from $(1,1)$ to $(n,m)$, where the weights are on the vertices. Here, the weight on the vertex $(i,j)$ is the distance $\left\|{p_{i}-q_{j}}\right\|$. In particular, a Fréchet morphing is an $x/y$-monotone path in $H$ from $(1,1)$ to $(n,m)$. The standard algorithm to do this, traverses the grid, say, by increasing rows $i$, and in each row by increasing column $j$, such that the value at $(i,j)$ is the maximum of the length of the leash of this configuration, together with the minimum solution for $(i-1,j)$ and $(i,j-1)$. This algorithm leads to a straightforward, $O(nm)$ time algorithm, for the discrete Fréchet distance, but the morphing computed might be inferior, see Figure 1.2 for such a bad example.
@@
For simplicity, assume that the pairwise distances between all pairs of points in the two sequences are unique. We would like to imagine that we have a retractable leash, that can become shorter at times, and the leash “resists” being longer than necessary. It is thus natural to ask for the morphing where the leash is somehow as short as possible at any point in time.
@@
Buchin et al. bbms-lcfm-19 refers to this Fréchet distance as locally correct, but we prefer the retractable labeling. The phrase retractable Fréchet was used by Buchin et al. bblmm-cfdrl-16 but in a different (and not formally defined) context than ours.
@@
One way to get less sensitivity for noise, is to compute the total area “swept” by the leash as the walk is being performed. In the discrete case, we simply add up the lengths of the leashes during the configurations in the walk. There is also work on extending this to the continuous setting mp-cdtwt-99, bbknp-klmct-20. For the continuous case this intuitively boils down to computing (or approximating) an integral along the morphing. See Section 4.5 for more details.
@@ 1.2.3 Algorithm engineering the Fréchet distance
More recently, Bringmann et al. bkn-wdfpa-21 presented an optimized implementation of the decider for the Fréchet distance, using heuristic pruning rules to avoid exploring unnecessary parts of the free-space diagram. Somewhat informally, Bringmann et al. bkn-wdfpa-21 build a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
@@ 1.3.1 Result I: A new algorithm for retractable discrete Fréchet
Maybe more importantly, in practice, one does not need to explore the whole space of $nm$ configurations (since we are in the discrete case, a configuration $(i,j)\in\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$ encodes the matching of $p_{i}$ with $q_{j}$), as the algorithm can stop as soon as it arrives at the destination configuration $(n,m)$. Informally, if the discrete Fréchet distance is “small” compared to the vast majority of pairwise distances (i.e., the two sequences are similar), then the algorithm only explores a small portion of the configuration space. Thus, this leads to an algorithm that is faster than the standard algorithm in many natural cases, while the output morphing is significantly better.
@@
Figure 1.3: Two curves, and their associated VE-Fréchet graph. Note, that every internal edge of the grid contains a portal (i.e., vertex of the graph), but in many cases the portals of two edges are co-located on their common vertex. For the edges adjacent to the starting and ending corners we set their portals to lie at the corners themselves.
@@ 1.3.2 Result II: A new distance and algorithm: VE-Fréchet
1.3.2.1 Traveling only through vertex-edge events.
@@
A somewhat similar idea was used by Munich and Perona mp-cdtwt-99, but they used it in the other direction – namely, in defining a better CDTW distance for two discrete sequences. However, it is already present (implicitly) in the original work of Alt and Godau ag-cfdbt-95 – indeed, their algorithm for the Weak Fréchet distance uses only the Vertex-Edge events (i.e., edges in the free space diagram). This boils down to solving the bottleneck shortest path problem in undirected graph. An in this case, this problem can be solved by computing the minimum spanning tree (e.g., by Prim’s algorithm. which is a variant of Dijkstra’s algorithm), as Alt and Godau do. For the directed case, one needs to use a variant of Dijkstra’s algorithm gt-atbop-88 – see Lemma 2.2. See also Buchin et al. bbdfj-ctmc-17 who also used a similar idea.
@@
1.3.3 Result III: New algorithm for computing the regular Fréchet distance
@@
1.3.3.1 The hunt for a monotone morphing.
@@
We refer to this process of adding vertices to the two curves as refinement. In practice, in many cases, one or two rounds of (carefully implemented) refinement are enough to isolate the maximum leash in the morphing from the non-monotonicity, and followed by the above brute-force monotonization leads to the (practically) optimal Fréchet distance. Even for pathological examples, after a few more rounds of refinement, this process computes the exact Fréchet distance (i.e., the computed lower bound, which is the VE-Fréchet distance, is equal to the width of the computed monotone morphing).
@@
As we are implementing our algorithm using floating point arithmetic, and the calculation of the optimal Fréchet distance involves distances, impreciseness is unavoidable. A slight improvement in preciseness can be achieved by using squared distances (and also slightly faster code). In particular, we take the somewhat pragmatic view, that an approximation to the optimal up to a factor of (say) $1.00001$ can be considered as computing the “optimal” solution.
1.3.4 Result IV: Computing the regular Fréchet distance quickly for real inputs
The above leaves us with a natural strategy for computing the Fréchet distance between two given curves. Compute quickly, using simplification, a morphing between the two input curves, and maintain (using VE-Fréchet, for example) both upper bound and lower bounds on the true Fréchet distance. By carefully inspecting the morphing, (re)simplifying the curves in a way that is sensitive to their (local) Fréchet distance, and recomputing the above bounds, one can get an improved morphing. Repeat this process potentially several times till the upper and lower bounds meet, at which point the optimal Fréchet distance has been computed.
This seems somewhat of an overkill but it enables us to compute (in practice) the exact Fréchet distance between huge polygonal curves quickly.
@@ 1.3.4 Result IV: Computing the regular Fréchet distance quickly for real inputs
1.3.5 Result V: Implementation in Julia and Python
We implemented the above algorithms as official packages in Python and Julia. A webpage with animations showing our algorithm in action is available here. The Python package is available at https://github.com/eliotwrobson/FrechetLib, and the Julia package is available at https://github.com/sarielhp/FrechetDist.jl. Animations and examples computed by the new algorithm are available at https://sarielhp.org/p/24/frechet_ve/examples/.
1.3.6 Miscellaneous results

### v3

@@ Abstract
We show that a variant of the continuous Fréchet distance between polygonal curves can be computed using essentially the same algorithm used to solve the discrete version. The new variant is not necessarily monotone, but this shortcoming can be easily handled via refinement.
Combined with a Dijkstra/Prim type algorithm, this leads to a realization of the Fréchet distance (i.e., a morphing) that is locally optimal (aka locally correct), that is both easy to compute, and in practice, takes near linear time on many inputs. The new morphing has the property that the leash is always as short-as-possible. These matchings/morphings are more natural, and are better than the ones computed by standard algorithms – in particular, they handle noise more graciously. This should make the Fréchet distance more useful for real world applications.
We implemented the new algorithm, and various strategies to obtain fast practical performance. We performed extensive experiments on our new algorithm, and released publicly available (and easily installable and usable) Julia and Python packages. In particular, the Julia implementation, for computing the regular Fréchet distance, seems to be significantly faster than other currently available implementations. See Figure 2.1.
Links and information about our implementations are available here: frechet.xyz, with numerous detailed examples with animations are available here.
@@ 1.1 Definitions
Given two polygonal curves, their Fréchet distance is the length of a leash that a person needs, if they walk along one of the curves, while a dog connected by the leash walks along the other curve, assuming they synchronize their walks so as to minimize the length of this leash. (I.e. they walk so as to minimize their maximum distance apart during the walk.) Our approach is slightly different than the standard approach, and we define it carefully first.
@@ 1.1.2 Fréchet distance
Conceptually, the Fréchet distance is the problem of computing the minimum bottleneck matching between two curves, respecting the order and continuity of the curves. [^2] Formally, since the reparameterization is not one-to-one, this is not quite a matching. One can restrict to using only such bijections, with no adverse effects, but it adds a level of tediousness, which we avoid for the sake of simplicity of exposition. Alternatively, it is an $L_{\infty}$-norm type measure of the similarity between two curves. It thus suffers from sensitivity to outliers. Furthermore, even if only a small portion of the morphing requires a long leash, the measure, and the algorithms computing it, may use this long leash in large portions of the walk, generating a matching that is loose in many places, see Figure 1.2.
@@
Figure 1.1: Two curves, their free space diagram (and the associated elevation function), and the optimal Fréchet morphing between the two curves encoded as an $x/y$-monotone curve. More illustrations and animations of this example are available here.
The classical (discrete) Fréchet morphing, caring only about the maximum leash length.
The retractable discrete Fréchet morphing, using the shortest leash possible at each point.
@@ 1.2 Background
Unfortunately, it is believed this problem requires quadratic time in the worst case, although a logarithmic speedup is possible; see bbmm-fswdi-17 and references therein. The quadratic time can be improved for realistic inputs by assuming that the input is “nice”, and introducing approximation, but the resulting algorithms are still not simple dhw-afdrc-12.
@@ 1.2.1 Variants of the Fréchet distance
The weak Fréchet distance allows morphings where the agents are allowed to go back and traverse portions of the curve visited previously (i.e., the morphing does not have to be $x/y$-monotone). Since the weak Fréchet distance allows considering more parameterizations, it is potentially smaller, and we have that $\mathsf{d}^{w}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, for any two curves $\pi,\sigma$. Elegantly, Alt and Godau ag-cfdbt-95 showed the weak Fréchet distance can be reduced to computing the minimum spanning tree of an appropriate graph. Unfortunately, there does not seem to be a natural way to overcome this non-monotonicity (and thus get the “strong” version).
@@
The complexity of these algorithms, together with sensitivity of the Fréchet distance to noise, lead to using “easier” related measures, such as the discrete version of the problem, and dynamic time-warping (discussed below). In the discrete version, you are given two sequences of points $p_{1},\ldots,p_{n}$, and $q_{1},\ldots,q_{m}$, and the purpose is for two “frogs” starting at $p_{1}$ and $q_{1}$, respectively, to jump through the points in the sequence until reaching $p_{n},q_{m}$, respectively, while minimizing the maximum distance between the two frogs during this traversal. At each step, only one of the frogs can jump from its current location to the next point in its sequence (no jumping back). Computing the optimal distance under this measure can be done by dynamic-programming, similar to the standard approach to edit-distance. Indeed, the configuration space here is the grid $H=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$, where $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$.
@@
For the (monotone) discrete Fréchet distance the induced graph on the grid $H$ is a DAG, and the task of computing the Fréchet distance is to find a minimum bottleneck path from $(1,1)$ to $(n,m)$, where the weights are on the vertices. Here, the weight on the vertex $(i,j)$ is the distance $\left\|{p_{i}-q_{j}}\right\|$. In particular, a Fréchet morphing is an $x/y$-monotone path in $H$ from $(1,1)$ to $(n,m)$. The standard algorithm to do this traverses the grid, say, by increasing rows $i$, and in each row by increasing column $j$, such that the value at $(i,j)$ is the maximum of the length of the leash of this configuration, together with the minimum solution for $(i-1,j)$ and $(i,j-1)$. This algorithm leads to a straightforward, $O(nm)$ time algorithm for the discrete Fréchet distance. However, the morphing computed might be inferior, see Figure 1.2 for such a bad example.
@@
For simplicity, assume that the pairwise distances between all pairs of points in the two sequences are unique. We would like to imagine that we have a retractable leash, that can become shorter at times, and the leash “resists” being longer than necessary. It is thus natural to ask for a morphing where the leash as short as possible at any point in time.
@@
Buchin et al. bbms-lcfm-19 refers to this Fréchet distance as locally correct, but we prefer the retractable labeling. The term “retractable Fréchet” was used by Buchin et al. bblmm-cfdrl-16, but in a different (and not formally defined) context than ours.
@@
One way to get less sensitivity for noise is to compute the total area “swept” by the leash as the walk is being performed. In the discrete case, we simply add up the lengths of the leashes during the configurations in the walk. There is also work on extending this to the continuous setting mp-cdtwt-99, bbknp-klmct-20. For the continuous case this intuitively boils down to computing (or approximating) an integral along the morphing. See Section 5.5 for more details.
@@ 1.2.3 Algorithm engineering the Fréchet distance
More recently, Bringmann et al. bkn-wdfpa-21 presented an optimized implementation of the decider for the Fréchet distance. Somewhat informally, Bringmann et al. bkn-wdfpa-21 builds a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
@@ 1.3.1 Result I: A new algorithm for retractable discrete Fréchet
Maybe more importantly, in practice, one does not need to explore the whole space of $nm$ configurations (since we are in the discrete case, a configuration $(i,j)\in\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$ encodes the matching of $p_{i}$ with $q_{j}$), as the algorithm can stop as soon as it arrives at the destination configuration $(n,m)$. Informally, if the discrete Fréchet distance is “small” compared to the vast majority of pairwise distances (i.e., the two sequences are similar), then the algorithm only explores a small portion of the configuration space. Thus, this leads to an algorithm that is faster than the standard algorithm in many natural cases, while computing a significantly better output morphing.
@@
Figure 1.3: Two curves, and their associated VE-Fréchet graph. Note, that every internal edge of the grid contains a portal (i.e., vertex of the graph), but in many cases the portals of two edges are co-located on their common vertex. For the edges adjacent to the starting and ending corners we set their portals to lie at the corners themselves. See here for more info.
@@ 1.3.2 Result II: A new distance and algorithm: VE-Fréchet
Traveling only through vertex-edge events.
@@
A somewhat similar idea was used by Munich and Perona mp-cdtwt-99, but they used it in the other direction – namely, in defining a better CDTW distance for two discrete sequences. However, this idea was already present (implicitly) in the original work of Alt and Godau ag-cfdbt-95 – indeed, their algorithm for the Weak Fréchet distance uses only the Vertex-Edge events (i.e., edges in the free space diagram). This boils down to solving the bottleneck shortest path problem in an undirected graph. In this case, this problem can be solved by computing the minimum spanning tree (e.g., by Prim’s algorithm, which is a variant of Dijkstra’s algorithm), as Alt and Godau do. For the directed case, one needs to use a variant of Dijkstra’s algorithm gt-atbop-88 – see Lemma 3.2. See also Buchin et al. bbdfj-ctmc-17 who also used a similar idea.
@@
1.3.3 Result III: New algorithm for the regular Fréchet distance
@@
The hunt for a monotone morphing.
@@
We refer to this process of adding vertices to the two curves as refinement. In practice, in many cases, one or two rounds of (carefully implemented) refinement are enough to isolate the maximum leash in the morphing from the non-monotonicity, and followed by the above brute-force monotonization leads to the (practically) optimal Fréchet distance. Even for pathological examples, after a few more rounds of refinement, this process computes the exact Fréchet distance. That is, the computed lower bound, which is the VE-Fréchet distance, is equal to the width of the computed monotone morphing, which is the Fréchet distance.
@@
As we are implementing our algorithm using floating point arithmetic, and the calculation of the optimal Fréchet distance involves distances, impreciseness is unavoidable. A slight improvement in preciseness can be achieved by using squared distances (and also slightly faster code) — but for simplicity we have not used this idea in our code. In particular, we take the somewhat pragmatic view, that an approximation to the optimal up to a factor of (say) $1.00001$ can be considered as computing the “optimal” solution.
Note that Fréchet morphings are somewhat less sensitive to numerical issues than other geometric problems — indeed, once a morphing is computed, one can compute its width directly.
1.3.4 Result IV: Computing the Fréchet distance quickly for real inputs
The above leaves us with a natural strategy for computing the Fréchet distance between two given curves. Compute quickly, using simplification, a morphing between the two input curves, and maintain (using VE-Fréchet, for example) both upper and lower bounds on the true Fréchet distance. By carefully inspecting the morphing, (re)simplifying the curves in a way that is sensitive to their (local) Fréchet distance, and recomputing the above bounds, one can get an improved morphing. Repeat this process potentially several times till the upper and lower bounds meet, at which point the optimal Fréchet distance has been computed.
This seems somewhat overkill, but it enables us to compute (in practice) the exact Fréchet distance between huge polygonal curves quickly.
@@ 1.3.4 Result IV: Computing the Fréchet distance quickly for real inputs
1.3.5 Main contribution: Implementation in Julia and Python
We implemented the above algorithms as official packages in Python and Julia. A webpage with animations showing our algorithm in action is available here. The Python package is available at https://github.com/eliotwrobson/FrechetLib, and the Julia package is available at https://github.com/sarielhp/FrechetDist.jl. Animations and examples computed by the new algorithm are available at https://frechet.xyz/.
1.3.6 Additional results